OpenStudy (shamil98):

What is the sum of the first five terms of a geometric series with a_1 = 6 and r = 1/3?

OpenStudy (shamil98):

Using the formula.. \[\sum_{k=0}^{n-1} (ar^k) = a \frac{ 1-r^n }{ 1-r }\]

OpenStudy (shamil98):

so.. \[6 \times \frac{ 1-(1/3)^5 }{ 1-1/3 }\] is this set up correctly?

OpenStudy (shamil98):

\[6 \times \frac{ 242/243 }{ 2/3 }\] \[6 \times \frac{ 242 }{ 243 } \times \frac{ 3 }{ 2 }\]

OpenStudy (shamil98):

\[6 \times \frac{ 121 }{ 81 }\]

OpenStudy (shamil98):

\[\frac{ 726 }{ 81 }\] is this the answer..